OpenStudy (anonymous):

7. Simplify the following product of complex numbers. (-1 + 5i) · (4 – 8i)

OpenStudy (anonymous):

(a+bi)(c+di) = (ac-bd) +(bc+ad)i

OpenStudy (anonymous):

(-1*4-5*8)+(5*8+-1*4)i

OpenStudy (anonymous):

mind the negatives (-8)

OpenStudy (anonymous):

so would it look like this (-1 + 5i)(4 - 8i) = -4 + 8i + 20i - 40i^2 = -4 + 28i + 40 = 36 + 28i

OpenStudy (anonymous):

yeah straight foil works perfect